In a sample of 50 people, 21 had type O blood, 22 had type A blood, 5 had type B blood, and
2 had type AB blood. Set up a frequency distribution and find the following probabilities.
a. A person has type O blood.
b. A person has type A or type B blood.
c. A person has neither type A nor type O blood.
d. A person does not have type AB blood.
1
Expert's answer
2021-08-03T16:31:23-0400

Frequency distribution:

A – 22

B – 5

AB – 2

O – 21

Total 50

a. "P(O) = \frac{21}{50}=0.42"

b. "P(A \;or\; B) = \frac{22+5}{50}=\frac{27}{50}=0.54"

c. "P(neither \;A \;nor\; O) = \frac{5+2}{50}= \frac{7}{50}= 0.14"

d. "P(not \;AB) = \frac{22+5+21}{50}= \frac{48}{50}=0.96"